OpenStudy (anonymous):

Help me please Trains A and B are traveling in same direction on parallel tracks. Train A is traveling at 80mph, and train B is traveling at 84 mph. Train A passes a station at 2:10am. If train B passes the same station at 2:40am what time will train B catch up with train A?

OpenStudy (stacey):

For this problem we use the problem rate * time = distance. If t represents the time that train B travels from the station until they catch up, then train A will have a time of (t+0.5) to travel that same distance.

OpenStudy (stacey):

For train A: distance = 80 * (t + 0.5) For train B: distance = 84 * t Since the distance from the station they pass to where they catch up is the same distance, we can set the two expressions equal to each other. 80(t + 0.5) = 84t

OpenStudy (anonymous):

-10?

OpenStudy (stacey):

You should get a positive 10. t = 10

OpenStudy (stacey):

80t + 40 = 84t 40 = 4t 10 = t

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

so 10am?

OpenStudy (stacey):

Since t represents the time of train B, we add that time to 2:40am to get the time that the train B catches up to train A.

OpenStudy (stacey):

2:40am + 10 hours = 12:40pm

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

oh, I see